# Gauge-Invariant Entire-Function Regulators and UV Finiteness in Non-Local Quantum Field Theory

### v3

@@ VIII Discussion and outlook
We have given a detailed and fully covariant derivation of how gauge-invariant entire-function regulators operate in HUFT and related nonlocal QFTs. The essential points can be summarized as follows: the regulator is implemented as an entire function $F(\Box/M_{*}^{2})$ of the covariant Laplace–Beltrami operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, where $D_{\mu}$ is the gauge- and diffeomorphism-covariant derivative and $M_{*}$ is the nonlocality scale. In the perturbative vacuum, where $g_{\mu\nu}\to\eta_{\mu\nu}$ and $A_{\mu}\to 0$, the operator $\Box$ reduces to the Minkowski d’Alembertian $\Box_{M}$, which is diagonalized by plane waves with eigenvalues $p^{2}$. Acting on a plane wave, the entire operator $F(\Box_{M}/M_{*}^{2})$ reduces to multiplication by $F(-p^{2}/M_{*}^{2})$. After Wick rotation, this becomes $F(-p_{E}^{2}/M_{*}^{2})$ in Euclidean momentum space, providing exponential UV damping of loop integrals for appropriate choices of $F$. Gauge invariance and BRST symmetry are preserved by constructing the regulator from background-covariant operators in the background-field formalism. The Ward and Slavnov–Taylor identities retain their standard form, and no unphysical degrees of freedom are introduced. Liouville’s theorem implies that nonconstant entire functions are unbounded and possess an essential singularity at infinity, but this has no physical consequences. Only the finite complex $p^{2}$-plane and the real Euclidean axis are probed by perturbative amplitudes, where $F(-p_{E}^{2}/M_{*}^{2})$ is analytic and free of zeros. The nonlocality associated with $F(\Box/M_{*}^{2})$ corresponds to a smearing of fields over a length scale $\ell_{*}\sim M_{*}^{-1}$. Locality and microcausality are recovered in the limit $M_{*}\to\infty$, and for large but finite $M_{*}$ the theory is quasi-local with negligible deviations at accessible scales.
@@
In the specific context of HUFT, these results provide a rigorous foundation for the use of entire-function regulators to render both gauge and gravitational sectors finite while maintaining a well-defined particle spectrum and preserving the geometric structure of the theory. They also clarify why Liouville’s theorem, often cited as an objection at a heuristic level, does not pose an obstacle to the physical viability of such regulators.

### v5

@@ IX Discussion and outlook
We have given a detailed and fully covariant derivation of how gauge-invariant entire-function regulators operate in non-local QFTs and related nonlocal QFTs. The essential points can be summarized as follows: the regulator is implemented as an entire function $F(\Box/M_{*}^{2})$ of the covariant Laplace–Beltrami operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, where $D_{\mu}$ is the gauge- and diffeomorphism-covariant derivative and $M_{*}$ is the nonlocality scale. In the perturbative vacuum, where $g_{\mu\nu}\to\eta_{\mu\nu}$ and $A_{\mu}\to 0$, the operator $\Box$ reduces to the Minkowski d’Alembertian $\Box_{M}$, which is diagonalized by plane waves with eigenvalues $p^{2}$. Acting on a plane wave, the entire operator $F(\Box_{M}/M_{*}^{2})$ reduces to multiplication by $F(-p^{2}/M_{*}^{2})$. After Wick rotation, this becomes $F(-p_{E}^{2}/M_{*}^{2})$ in Euclidean momentum space, providing exponential UV damping of loop integrals for appropriate choices of $F$. Gauge invariance and BRST symmetry are preserved by constructing the regulator from background-covariant operators in the background-field formalism. The Ward and Slavnov–Taylor identities retain their standard form, and no unphysical degrees of freedom are introduced. Liouville’s theorem implies that nonconstant entire functions are unbounded and possess an essential singularity at infinity, but this has no physical consequences. Only the finite complex $p^{2}$-plane and the real Euclidean axis are probed by perturbative amplitudes, where $F(-p_{E}^{2}/M_{*}^{2})$ is analytic and free of zeros. The nonlocality associated with $F(\Box/M_{*}^{2})$ corresponds to a smearing of fields over a length scale $\ell_{*}\sim M_{*}^{-1}$. Locality and microcausality are recovered in the limit $M_{*}\to\infty$, and for large but finite $M_{*}$ the theory is quasi-local with negligible deviations at accessible scales.